Find three consecutive even integers such that the sum of the least integer and the greatest integer is 12 more than the middle integer
Answer by richwmiller(17219)

a+a+4=12+a+2
2a+4=a+14
2a=a+10
a=10
10,12, 14
check
10+14=(12)+12
24=24
ok